The following information relates to the debt securities investments of Sheffield Company. 1. On February 1, the company purc(3) Jul. 1 (4) Sep. 1 (To record interest.) (To record adjustment.)
Answer #1

Solution:

No.of Event Date Account Title and Explanation Debit ($) Credit ($)
1. 1 Feb. Investment in Bond - Gibbons co. 318,000
Interest receivable (318,000 × 11% ×4/12) 11,660
Cash 329,660
(To record investment in bond)
2. 1Apr. Cash 17,490
Interest receivable 11,660
Interest revenue(318,000 × 11% × 2/12) 5,830
(To record semiannual interest received)
3. 1 July. Investments in Bonds - Sampson Inc. 219,600
Interest receivable (219,600 × 9% × 1/12) 1,647
Cash 221,247
(To record purchase bond)
4. 1 Sep. Cash 56,074
Available for sale securities 55,200
Realized gain on the sale of available for sale securities 874
(To record realized gain)
5. 1.Oct. Cash (318,000 × 11% × 6/12) 17,490
Interest income 17,490
(To record interest income)
6. 1Dec. Cash (219,600 × 9% ×5/12) 8,235
Interest income 8,235
( To record interest income)
7. Gibbons co.
31Dec. Available for sale securities ($318,000 ×93% =$295,740)-$318,000 =$22,260 22,260
Unrealized loss other comprehensive income 22,260
Sampson inc.
Available for sale securities ($219,200 ×91% -199,472) -$219,200 = $19,728 19,728
Unrealized loss other comprehensive income 19,728
( To record adjustment)

Explanation:

Sep. 1 = ($55,200 × 97% =$53,544) +(55200 × 11% ×5/12 =$2,530) =$56,074 - $55,200 =$874